OpenStudy (anonymous):

If the graph of the line x-2y=-1 in the xy plane is reflected across the x axis, which of the following is an equation of the resulting reflection? x-2y=-1 x+2y=-1 x-2y=1 x+2y=1 2x+y=-2 Show work!

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Directrix lol.

OpenStudy (debbieg):

For y=f(x), the equation y=-f(x) is a reflection over the x-axis. That is, it takes every y coordinate and multiplies it by -1. So if you just solve the equation for y=mx+b, then the reflection across the x-axis is the line y=-(mx+b). Then you just need to put it back into standard form to make it fit your answer choices.